OpenStudy (22chocolate22):

What is the square root of this dang life divided by a hopeless answers, times dreams that stink? hehe

OpenStudy (christopher1448):

??????

OpenStudy (joshmorris):

hi

OpenStudy (22chocolate22):

Hello?

OpenStudy (bonnieisflash1.0):

welcome to five nights at freddy

OpenStudy (shadowlegendx):

\[\frac{ \sqrt dang + life }{ hopeless + answers } \times dreams+ that+ stink\]

OpenStudy (22chocolate22):

Thanks ShadowLegandX. Thanks.

OpenStudy (shadowlegendx):

No problem

OpenStudy (22chocolate22):

I'm fanning because of that.

OpenStudy (christopher1448):

FAN ME PLZ